OpenStudy (anonymous):

find the equation of the line containing the given pair of points. (1,5),(3,6) y= ?

OpenStudy (anonymous):

y-y1 = ((y2-y1)/(x2-x1))(x-x1)

OpenStudy (abmon98):

6-5/3-1=1/2 thats the gradient apply the law y-y1=m(x-x1) y-5=1/2(x-1) 2y-10=x-1 y=1/2x+9